# EKF–SINDy: Empowering the extended Kalman filter with sparse identification of nonlinear dynamics

### v1

@@ Abstract
Observed data from a dynamic system can be assimilated into a predictive model by means of Kalman filters. Nonlinear extensions of the Kalman filter, such as the Extended Kalman Filter (EKF), are required to enable the joint estimation of (possibly nonlinear) system dynamics and of input parameters. To construct the evolution model used in the prediction phase of the EKF, we propose to rely on the Sparse Identification of Nonlinear Dynamics (SINDy). The numerical integration of a SINDy model leads to great computational savings compared to alternate strategies based on, e.g., finite elements. Indeed, SINDy allows for the immediate definition of the Jacobian matrices required by the EKF to identify system dynamics and properties, a derivation that is usually extremely involved with physical models. As a result, combining the EKF with SINDy provides a computationally efficient, easy-to-apply approach for the identification of nonlinear systems, capable of robust operation even outside the range of training of SINDy. To demonstrate the potential of the approach, we address the identification of a linear non-autonomous system consisting of a shear building model excited by real seismograms, and the identification of a partially observed nonlinear system. The challenge arising from applying SINDy when the system state is not accessible has been relieved by means of time-delay embedding. The great accuracy and the small uncertainty associated with the state identification, where the state has been augmented to include system properties, underscores the great potential of the proposed strategy, paving the way for the development of predictive digital twins in different fields.
@@ 1 Introduction
Dynamic system identification is crucial to enable the construction of predictive digital twins art:Wagg20 , and to implement control strategies for engineering systems with applications to condition-based maintenance of civil structures art:Matt24 , or to fuel consumption reduction in air transportation art:Guastoni23 . Despite advances in Machine Learning (ML) and Data Science art:Jordan15 , identifying explainable, reduced dimensional models of dynamic processes from big data is an open field of research art:Brunton16 . Searching for the physical relations that underlie a certain dynamic process is probably the best way to obtain robust and generalisable models, a characteristic uncommon in most of ML techniques. Consequently, making reliable predictions in scenarios lacking collected data is generally challenging or impossible art:Olivier21 . This type of uncertainty is commonly referred to as epistemic uncertainty. To address this challenge, we employ the Sparse Identification of Nonlinear Dynamics (SINDy) proposed in art:Brunton16 . SINDy constructs robust and generalisable models by assuming that only a few important terms govern the dynamics of the considered system. This assumption holds for many physical systems when an appropriate basis is used to describe the space of functions governing their dynamics art:Brunton16 .
@@
To construct a digital twin, the model must evolve over time to reflect potential changes in its physical counterpart. This evolution should be, possibly online, driven by real-world data art:Wagg24 . In this work, we have exploited an Extended Kalman filters (EKF) to perform data assimilation art:Kalman60 leveraging a prediction-correction scheme. During the prediction stage, the SINDy model evolves the system state. The procedure can also accommodate potential external forcing rendering the system not autonomous. During the correction stage, acquired data are used to refine previous estimates, update the physical parameters governing the system dynamics, and potentially reduce the associated uncertainties. As a result, a joint estimation of the system is achieved (termed joint because both the state of the system and the physical parameters underlying the dynamics are jointly updated, see proc:Wan00 ). A schematic representation of the EKF-SINDy procedure is reported in Fig. 1. More details on this methodology will be discussed in Sec. 2.
@@
In realistic applications, the observation variables assimilated by the EKF do not necessarily match up with the system state on which SINDy is performed. This mismatch arises in two scenarios: where the observations are fewer than the system state variables, potentially limiting the full description of the system dynamics, and where observations are in excess, potentially overwhelming the EKF-SINDy framework with redundant information. In the former case, we can resort to uplifting techniques such as, e.g., embedding techniques art:Brunton17 ; bakarji2023discovering or recurrent decoder networks williams2023sensing , to recover the hidden state components from the available, observed time-series. To address the latter case, dimensionality reduction techniques, employing proper orthogonal decomposition and/or encoder neural networks, have been coupled with SINDy to simultaneously reduce dimensionality while learning the corresponding governing dynamical system as in champion2019data ; bakarji2023discovering ; conti2023reduced . In this work, we deal with the first scenario, focusing on revealing hidden variables when only partial observations are available. The integration in the method of dimensionality reduction techniques to deal with high-dimensional data will be the focus of future work.
@@
Figure 1: EKF-SINDy methodology. In the prediction phase, the dynamical system identified by SINDy advances the augmented state $\boldsymbol{\varkappa}=[\mathbf{x}^{\top},\boldsymbol{\phi}^{\top}]^{\top}$—comprising system state $\mathbf{x}$ and parameters $\boldsymbol{\phi}$—to a prior estimate $\boldsymbol{\varkappa}^{-}$; the correction phase involves the EKF assimilating system observations $\bar{\mathbf{y}}$, resulting in the updated augmented state $\boldsymbol{\varkappa}^{+}$.
@@
In the recent literature, other works employ a ML-based identified model in the prediction phase of the KF. In proc:Coskun , predictions were made using recurrent Neural Networks (NN), specifically Long Short-Term Memory (LSTM) modules; LSTM modules were also utilised to model the process and observation noise required by the filter. The linear KF was used to estimate only the dynamics of the observed system. In contrast, our work explicitly calibrates the physical parameters affecting the dynamical process, like in art:Liu24 where NNs and EKF were combined to perform joint estimation of mechanical systems. This work improved upon previously proposed variational autoencoders-based approaches, such as the deep Markov models proc:Krishnan17 , by relying on the EKF equation to infer the latent variables governing the system dynamics. Compared to art:Liu24 , our methodology looks better suited for a physical process whose dynamics is controlled by a few important terms. This is because NNs require large volume of training data, do not easily account for known symmetries and constraints, and struggle to handle epistemic uncertainties art:Kaiser18 .
@@
SINDy has previously been coupled with the KF in proc:Wang22 . In that work, the estimation procedure directly targeted the SINDy coefficients. However, as previously mentioned, relying on the linear KF precludes the possibility of performing joint estimation. At variance with proc:Coskun , the proposal of proc:Wang22 focused on identifying the coefficients of dynamic models without correcting the state predictions.
@@
The application of the proposed approach to experimental data appears to be extremely promising. SINDy can be directly identified from preliminary acquired data. In this study, we have tested our methodology using simulated data for two primary reasons: first, to showcase the method potential and its ability to handle non autonomous and partially observed nonlinear systems; moreover, to demonstrate that the proposed methodology may be advantageous even when a physics-based model of the system is available. In the latter case, the model identified by SINDy serves as a surrogate for the physics-based model enabling significant computational time savings compared to the integration of nonlinear models. Relying on fast computational approaches has recently proven crucial in many applications, such as the design optimisation of Micro-Electro Mechanical Systems (MEMS) art:Vizzaccaro24 , and inverse analysis in structural health monitoring art:CAS21 ; chp:CSAI22 , just to mention two relevant cases dealing with structural dynamics, both at the micro- and the macro-scale. The need of computational efficiency also underlies the use of SINDy in control applications art:Kaiser18 ; art:abdullah23 . Alternatively, a recent work proposed to reduce the computational burden of the estimation process by using a rank-reduced version of the KF proc:Schmidt23 . Another major advantage of using SINDy is the greatly simplified computation of the Jacobian matrices required by the EKF formulation, compared to the involved derivation typically encountered even for relatively simple mechanical systems art:RUENG24 .
@@
The remainder of the paper is organized as follows. The methodology is presented in Sec. 2: first, KF is introduced; then the rationale behind SINDy is discussed; finally, the application of EKF-SINDy to joint estimation of autonomous and non-autonomous dynamic systems is detailed. Two case studies are presented in Sec. 3. In Sec. 3.1, a first case discusses the identification of a shear building subjected to a seismic event. Real seismograms are used as excitations, demonstrating the capability of our procedure of handling non-autonomous systems. A second case study, presented in Sec. 3.2, focuses on the identification of a partially observed nonlinear resonator. To address the impossibility to observe the whole system state, we precede the training of the SINDy model by applying the time-delay embedding art:Champion19 . Final considerations are collected in Sec. 4, along with a discussion of future developments. The source code of the proposed method is made available in the public repository EKF-SINDy EKFSINDY_repo .
@@ 2.1 Extended Kalman Filtering
Considering a dynamic system of interest, Kalman filters exploit a state–space representation of the type:
@@
where: $\mathbf{x}(t)\in\mathbb{R}^{n}$ is the state vector of the system at time $t$; $\mathbf{f}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is the function of $\mathbf{x}(t)$ describing the dynamics of the system. Nonlinear versions of the KF, like the EKF, are required if $\mathbf{f}$ is nonlinear. Hereon, we rely on the extended Kalman filter (EKF); the reader may refer, e.g., to book:Simon06_13 for a complete derivation and analysis of the EKF theory.
Dynamic systems can be used to model the response of a building to seismic excitations, or the behavior of MEMS book:Corigliano4 . In general, $\mathbf{f}$ is not exactly known. Thus, it becomes important to assimilate incoming data of the system to update the model and the resulting predictions. For instance, these data can consist of floor acceleration measurements for a building art:Farrar98 , or displacement measurements for capacitive sensors. According to the discrete acquisition of system measurements, data assimilation is performed at discrete $t_{j}$ instants of observation, with $j=1,\ldots,T$. As state variables may not be directly compared against incoming data, the following observation equation:
$\mathbf{y}(t_{j})=\mathbf{h}(\mathbf{x}(t_{j})),$ (2)
s usually introduced to encode the state-to-observation map. Specifically: $\mathbf{h}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{o}$ is a possibly nonlinear function; $\mathbf{y}(t_{j})\in\mathbb{R}^{o}$ are the quantities to be compared with the incoming observations $\bar{\mathbf{y}}(t_{j})\in\mathbb{R}^{o}$. According to Eqs. (1) and (2), model predictions are performed employing a continuous time description, while data assimilation is carried out at discrete time steps. For this reason, a hybrid (continuous–discrete) formulation of the KF will be considered in the following.
A prediction–correction scheme is then adopted to perform data assimilation. In the prediction phase, the state of the system is advanced in time through $\mathbf{f}$. This prediction is then corrected online through the comparison with incoming data. Hereon, we will use the superscript “--” to refer to quantities not yet updated by the correction stage, and the superscript “$+$” for quantities updated by the correction stage.
@@
where: $\mathbf{q}_{x}$ is the process noise vector, assumed to be sampled from a white, zero mean stochastic process $\mathbf{q}_{x}\sim\mathcal{N}(\mathbf{0},\mathbf{Q}_{x})$ with diagonal covariance matrix $\mathbf{Q}_{x}\in\mathbb{R}^{n\times n}$; $\mathbf{r}\in\mathbb{R}^{o}$ is the observation noise vector, assumed to be sampled from a white, zero-mean stochastic process $\mathbf{r}\sim\mathcal{N}(\mathbf{0},\mathbf{R})$ with diagonal covariance matrix $\mathbf{R}\in\mathbb{R}^{o\times o}$. The vector $\mathbf{q}_{x}$ accounts for uncertainties in the mapping $\mathbf{f}$, while $\mathbf{r}_{x}$ accounts for the noise affecting the incoming data. Their trade-off determines how much the filter relies on model predictions with respect to the acquired data book:Simon06_5 .
The Kalman filter follows the system evolution by looking at $\hat{\mathbf{x}}$ and $\mathbf{P}$, where: $\hat{\mathbf{x}}=\mathbb{E}[\mathbf{x}]$ is the mean value of $\mathbf{x}$; $\mathbb{E}$ computes the expected value of a quantity; $\mathbf{P}=\mathbb{E}[(\mathbf{x}-\hat{\mathbf{x}})(\mathbf{x}-\hat{\mathbf{x}})^{\top}]$ is the covariance matrix associated to $\mathbf{x}$.
@@
where $\mathbf{F}$ is the Jacobian of $\mathbf{f}(\mathbf{x}(t))$. The matrix $\mathbf{F}^{-}(\tau)$ is evaluated as:
@@
Different methods can be used to integrate Eq. (2.1). For its computational efficiency, we have considered the Euler forward method leading to:
@@
where $\mathbf{F}^{+}_{j-1}$ is evaluated for $\hat{\mathbf{x}}^{+}_{j-1}$ at $t_{j-1}$. In general, $\mathbf{Q}_{x}$ may modify in time; however, we neglect such possibility in this work. Moreover, we assume a constant time sampling $\Delta t=t_{j}-t_{j\!-\!1}$.
The Euler forward method is here a first-order explicit time integration technique, meaning that the local truncation error generated scales linearly with the time step $\Delta t$, provided this latter is chosen to be sufficiently small. More accurate methods can be employed, such as, e.g., higher-order Runge–Kutta methods. However, in the next section we will show that the local truncation error performed by the Euler forward method does not preclude good outcomes for the system estimation, also thanks to the adopted prediction–correction strategy. Similarly, implicit time stepping schemes can also be considered in order not to strongly constrain the choice of the time step, but it implies the solution of a (non)linear system of equations at each time step, in the case of (non)linear dynamic systems.

### v2

@@ Abstract
Measured data from a dynamical system can be assimilated into a predictive model by means of Kalman filters. Nonlinear extensions of the Kalman filter, such as the Extended Kalman Filter (EKF), are required to enable the joint estimation of (possibly nonlinear) system dynamics and of input parameters. To construct the evolution model used in the prediction phase of the EKF, we propose to rely on the Sparse Identification of Nonlinear Dynamics (SINDy). SINDy enables to identify the evolution model directly from preliminary acquired data, thus avoiding possible bias due to wrong assumptions and incorrect modelling of the system dynamics. Moreover, the numerical integration of a SINDy model leads to great computational savings compared to alternate strategies based on, e.g., finite elements. Last, SINDy allows an immediate definition of the Jacobian matrices required by the EKF to identify system dynamics and properties, a derivation that is usually extremely involved with physical models. As a result, combining the EKF with SINDy provides a data-driven computationally efficient, easy-to-apply approach for the identification of nonlinear systems, capable of robust operation even outside the range of training of SINDy. To demonstrate the potential of the approach, we address the identification of a linear non-autonomous system consisting of a shear building model excited by real seismograms, and the identification of a partially observed nonlinear system. The challenge arising from the use of SINDy when the system state is not entirely accessible has been relieved by means of time-delay embedding. The great accuracy and the small uncertainty associated with the state identification, where the state has been augmented to include system properties, underscores the great potential of the proposed strategy, paving the way for the setting of predictive digital twins in different fields.
@@ 1 Introduction
Dynamical system identification is crucial to enable the construction of predictive digital twins [1], and to implement control strategies for engineering systems with applications to condition-based maintenance of civil structures [2], or to fuel consumption reduction in air transportation [3]. Despite advances in Machine Learning (ML) and Data Science [4], identifying explainable, reduced dimensional models of dynamic processes from big data is an open field of research [5]. Searching for the physical relations that underlie a certain dynamic process is probably the best way to obtain robust and generalisable models, a characteristic uncommon in most of ML techniques. Consequently, making reliable predictions in scenarios lacking collected data is generally challenging or impossible [6]. This type of uncertainty is commonly referred to as epistemic uncertainty. To address this challenge, we employ the Sparse Identification of Nonlinear Dynamics (SINDy) proposed in [5]. SINDy constructs robust and generalisable models by assuming that only a few important terms govern the dynamics of the considered system. This assumption holds for many physical systems when an appropriate basis is used to describe the space of functions governing their dynamics [5].
@@
To construct a digital twin, the model must evolve over time to reflect potential changes in its physical counterpart. This evolution should be, possibly online, driven by real-world data [7]. In this work, we have exploited an Extended Kalman filters (EKF) to perform data assimilation [8] leveraging a prediction-correction scheme. During the prediction stage, the SINDy model evolves the system state. The procedure can also accommodate potential external forcing rendering the system not autonomous. During the correction stage, acquired data are used to refine previous estimates, update the physical parameters governing the system dynamics, and potentially reduce the associated uncertainties. As a result, a joint estimation of the system is achieved (termed joint because both the state of the system and the physical parameters underlying the dynamics are jointly updated, see [9]). A schematic representation of the EKF-SINDy procedure is reported in Fig. 1. More details on this methodology will be discussed in Sec. 2.
@@
In realistic applications, the observation variables assimilated by the EKF do not necessarily match up with the system state on which SINDy is performed. This mismatch arises in two scenarios: where the observations are fewer than the system state variables, potentially limiting the full description of the system dynamics, and where observations are in excess, potentially overwhelming the EKF-SINDy framework with redundant information. In the former case, we can resort to uplifting techniques such as, e.g., embedding techniques [10, 11] or recurrent decoder networks [12], to recover the hidden state components from the available, observed time-series. To address the latter case, dimensionality reduction techniques, employing proper orthogonal decomposition and/or encoder neural networks, have been coupled with SINDy to simultaneously reduce dimensionality while learning the corresponding governing dynamical system as in [13, 11, 14]. In this work, we deal with the first scenario, focusing on revealing hidden variables when only partial observations are available. The integration in the method of dimensionality reduction techniques to deal with high-dimensional data will be the focus of future investigations.
@@
Figure 1: EKF-SINDy methodology. In the prediction phase, the dynamical system identified by SINDy, together with random walk equation modelling the parameter evolution, advances the augmented state $\boldsymbol{\varkappa}=[\mathbf{x}^{\top},\boldsymbol{\phi}^{\top}]^{\top}$—comprising system state $\mathbf{x}$ and parameters $\boldsymbol{\phi}$—to a prior estimate $\boldsymbol{\varkappa}^{-}$; the correction phase involves the EKF assimilating system observations $\bar{\mathbf{y}}$, resulting in the updated augmented state $\boldsymbol{\varkappa}^{+}$.
@@
In the recent literature, other works employ a ML-based identified model in the prediction phase of the KF. In [15], predictions were made using recurrent Neural Networks (NN), specifically Long Short-Term Memory (LSTM) modules; LSTM modules were also utilised to model the process and observation noise required by the filter. The linear KF was used to estimate only the dynamics of the observed system. In contrast, our work explicitly calibrates the physical parameters affecting the dynamical process, like in [16] where NNs and EKF were combined to perform joint estimation of mechanical systems. This work improved upon previously proposed variational autoencoders-based approaches, such as the deep Markov models [17], by relying on the EKF equation to infer the latent variables governing the system dynamics. Compared to [16], our methodology looks better suited for a physical process whose dynamics is controlled by a few important terms. This is because NNs require large volume of training data, do not easily account for known symmetries and constraints, and struggle to handle epistemic uncertainties [18].
@@
SINDy has previously been coupled with the KF in [19]. In that work, the estimation procedure directly targeted the SINDy coefficients. However, as previously mentioned, relying on the linear KF precludes the possibility of performing joint estimation. At variance with [15], the proposal of [19] focused on identifying the coefficients of dynamic models without correcting the state predictions.
The application of the proposed approach to experimental data appears to be extremely promising. Indeed, SINDy can be directly applied to preliminary acquired data, thereby mitigating the risk of introducing a bias when modelling the system dynamics. If data driven model identification techniques like SINDy are not used, this bias can negatively impact the filtering outcome. In such cases, other possible mitigation strategies, for example based on the recently proposed integration with echo state networks, must be considered [20].
In this study, we have tested our methodology using simulated data for two primary reasons: first, to showcase the method potential and its ability to handle non autonomous and partially observed nonlinear systems; moreover, to demonstrate that the proposed methodology may be advantageous even when a physics-based model of the system is available. In the latter case, the model identified by SINDy serves as a surrogate for the physics-based model enabling significant computational time savings compared to the integration of nonlinear models. These computational savings are relevant to the online application of the procedure, in which data assimilation is performed by the EKF using a trained SINDy model to evolve the system dynamics. The training of the SINDy model, on the other hand, is carried out in an offline phase, which may be computationally intensive.
@@
Relying on fast computational approaches has recently proven crucial in many applications, such as the design optimisation of Micro-Electro Mechanical Systems (MEMS) [21], and inverse analysis in structural health monitoring [22, 23], just to mention two relevant cases dealing with structural dynamics, both at the micro- and the macro-scale. The need of computational efficiency also underlies the use of SINDy in control applications [18, 24]. Alternatively, a recent work proposed to reduce the computational burden of the estimation process by using a rank-reduced version of the KF [25]. Another major advantage of using SINDy is the greatly simplified computation of the Jacobian matrices required by the EKF formulation, compared to the involved derivation typically encountered even for relatively simple mechanical systems [26].
@@
The remainder of the paper is organized as follows. The methodology is presented in Sec. 2: first, KF is introduced; then the rationale behind SINDy is discussed; finally, the application of EKF-SINDy to joint estimation of autonomous and non-autonomous dynamical systems is detailed. Two case studies are presented in Sec. 3. In Sec. 3.1, a first case discusses the identification of a shear building subjected to a seismic event. Real seismograms are used as excitations, demonstrating the capability of our procedure of handling non-autonomous systems. A second case study, presented in Sec. 3.2, focuses on the identification of a partially observed nonlinear resonator. To address the impossibility to observe the whole system state, we precede the training of the SINDy model by applying the time-delay embedding [27]. Final considerations are collected in Sec. 4, along with a discussion of future developments. The source code of the proposed method is made available in the public repository EKF-SINDy [28].
@@ 2.1 Extended Kalman Filtering
Considering a dynamical system of interest, Kalman filters exploit a state–space representation of the type:
@@
where: $\mathbf{x}(t)\in\mathbb{R}^{n}$ is the state vector of the system at time $t$; $\mathbf{f}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is the function of $\mathbf{x}(t)$ describing the dynamics of the system. Nonlinear versions of the KF, like the EKF, are required if $\mathbf{f}$ is nonlinear. Hereon, we rely on the extended Kalman filter (EKF); the reader may refer, e.g., to [29] for a complete derivation and analysis of the EKF theory.
Dynamical systems can be used to model the response of a building to seismic excitations, or the behavior of MEMS [30]. In general, $\mathbf{f}$ is not exactly known. Thus, it becomes important to assimilate incoming data of the system to update the model and the resulting predictions. For instance, these data can consist of floor acceleration measurements for a building [31], or displacement measurements for capacitive sensors. According to the discrete acquisition of system measurements, data assimilation is performed at discrete $t_{j}$ instants of observation, with $j=1,\ldots,T$. As state variables may not be directly compared against incoming data, the following observation equation:
$\mathbf{y}(t_{j})=\mathbf{h}(\mathbf{x}(t_{j}))$ (2)
is usually introduced to encode the state-to-observation map. Specifically: $\mathbf{h}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{o}$ is a possibly nonlinear function; $\mathbf{y}(t_{j})\in\mathbb{R}^{o}$ are the quantities to be compared with the incoming observations $\bar{\mathbf{y}}(t_{j})\in\mathbb{R}^{o}$. According to Eqs. (1) and (2), model predictions are performed employing a continuous time description, while data assimilation is carried out at discrete time steps. For this reason, a hybrid (continuous–discrete) formulation of the KF will be considered in the following.
A prediction–correction scheme is then adopted to perform online data assimilation. In the prediction phase, the state of the system is advanced in time through $\mathbf{f}$. This prediction is then corrected through the comparison with incoming data. Hereon, we will use the superscript “--” to refer to quantities not yet updated by the correction stage, and the superscript “$+$” for quantities updated by the correction stage.
@@
where: $\mathbf{q}_{x}$ is the process noise vector, assumed to be sampled from a white, zero mean stochastic process $\mathbf{q}_{x}\sim\mathcal{N}(\mathbf{0},\mathbf{Q}_{x})$ with diagonal covariance matrix $\mathbf{Q}_{x}\in\mathbb{R}^{n\times n}$; $\mathbf{r}\in\mathbb{R}^{o}$ is the observation noise vector, assumed to be sampled from a white, zero-mean stochastic process $\mathbf{r}\sim\mathcal{N}(\mathbf{0},\mathbf{R})$ with diagonal covariance matrix $\mathbf{R}\in\mathbb{R}^{o\times o}$. The vector $\mathbf{q}_{x}$ accounts for uncertainties in the mapping $\mathbf{f}$, while $\mathbf{r}$ accounts for the noise affecting the incoming data. Their trade-off determines how much the filter relies on model predictions with respect to the acquired data [32].
The Kalman filter follows the system evolution by looking at $\hat{\mathbf{x}}$ and $\mathbf{P}$, where: $\hat{\mathbf{x}}=\mathbb{E}[\mathbf{x}]$ is the mean value of $\mathbf{x}$; $\mathbb{E}$ computes the expected value of a quantity; $\mathbf{P}=\mathbb{E}[(\mathbf{x}-\hat{\mathbf{x}})(\mathbf{x}-\hat{\mathbf{x}})^{\top}]$ is the covariance matrix associated to $\mathbf{x}$. This is equivalent to assume that $\mathbf{x}$ features a Gaussian distribution. This assumption is quite natural in system identification. A Gaussian probability distribution is associated with $\mathbf{x}$ and $\mathbf{y}$ because it is the distribution with maximum information entropy given the specified mean and variance [33].
@@
where $\mathbf{F}\in\mathbb{R}^{n\times n}$ is the Jacobian of $\mathbf{f}(\mathbf{x}(t))$. The matrix $\mathbf{F}^{-}(\tau)$ is evaluated as:
@@
Different methods can be used to integrate Eq. (4). For its computational efficiency, we have considered the Euler forward method leading to:
@@
where $\mathbf{F}^{+}(t_{j-1})$ is evaluated for $\hat{\mathbf{x}}^{+}_{j-1}$ at $t_{j-1}$. In general, $\mathbf{Q}_{x}$ may modify in time; however, we neglect such possibility in this work. Moreover, we assume a constant time step $\Delta t=t_{j}-t_{j\!-\!1}$.
The Euler forward method is a first-order explicit time integration technique, meaning that the local truncation error generated scales linearly with the time step $\Delta t$, provided this latter is chosen to be sufficiently small. This time integration method has been selected for its simplicity. However, more accurate methods can be employed, such as, e.g., higher-order Runge–Kutta methods, as the selection of the time integration technique does not affect the generality of the proposed procedure. In the next section we will show that the local truncation error performed by the Euler forward method does not preclude good outcomes for the system estimation, also thanks to the adopted prediction–correction strategy. Similarly, implicit time stepping schemes can also be considered in order not to strongly constrain the choice of the time step, but it implies the solution of a (non)linear system of equations at each time step, in the case of (non)linear dynamical systems. If the employed integration algorithm is not unconditionally stable, as typical of explicit time integration schemes, preliminary tests must be performed to verify that $\Delta t$ does not induce any loss of convergence in the integration of system dynamics.